• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name cusanno

    Log entry time 08:00:24 on March15,2006

    Entry number 166756

    keyword=Owl Shift Summary

    Shift begins in the middle of changes: new DC voltages, CODA crash, target on way of restoring.

    00:50 - we set and check the new DC voltages and restored DAQ (ROC23 reboot).
    Run 2979 is test of new DC HV setting, ps2=50, ps3=1

    01:45 - NMR measurements.
    Some troubles during measurement, we called Ameya.

    02:05 - Polarization is well above 30% (HALOG 166718), so we switch to production.
    Compton background is still not so good.

    02:30 - Justin noticed raster size smaller than 2 mm x 2 mm. Asked MCC to have larger raster.
    Now MCC is working with 4.5 x 4.5 and presently this give us
    2 mm x 2 mm at target position as desired.
    SLI shows energy spread around 1.3 x 10-4, MCC says they can not do much better for the night.
    May it be related to the Compton background in some way?

    04:00 - Noticed that raster is still slightly smaller than 2 mm in x direction, we asked MCC to have 5 x 5.

    04:29 - Beam away for three minutes due to cavity problems.

    05:25 - NMR measurement performed

    07:00 - Stop production, beam away for beam studies.
    No more beam at least until the afternoon.

    07:20 - Controlled Access

    Shift end in controlled access.


    production run:

    2980, 670 k events, ended to check raster size
    2981, 2M events
    2982, 2M events
    2983, 850k events. ended because beam away
    2984, 2 M events
    2985 - JUNK - NMR measurement needed
    2986, 2M events
    2987, 2M events
    2988, 106k events, ended because beam is away for beam studies